the area of a rectangular carpet is 18m squared. the length of the carpet is 3m longer than the width. how long is the carpet?

Answer by checkley79(3341)

AREA=LW
L=W+3
18=(W+3)W
18=W^2+3W
W^2+3W-18=0
(W+6)(W-3)=0
W-3=0
W=3 ANS. FOR THE WIDTH.
L=3+3=6 ANS. FOR THE LENGTH.
PROOF:
18=6*3
18=18